# End-to-End Chess Recognition

### v1

@@ Abstract
Chess recognition refers to the task of identifying the chess pieces configuration from a chessboard image. Contrary to the predominant approach that aims to solve this task through the pipeline of chessboard detection, square localization, and piece classification, we rely on the power of deep learning models and introduce two novel methodologies to circumvent this pipeline and directly predict the chessboard configuration from the entire image. In doing so, we avoid the inherent error accumulation of the sequential approaches and the need for intermediate annotations.
Furthermore, we introduce a new dataset, Chess Recognition Dataset (ChessReD), specifically designed for chess recognition that consists of 10,800 images and their corresponding annotations. In contrast to existing synthetic datasets with limited angles, this dataset comprises a diverse collection of real images of chess formations captured from various angles using smartphone cameras; a sensor choice made to ensure real-world applicability. We use this dataset to both train our model and evaluate and compare its performance to that of the current state-of-the-art. Our approach in chess recognition on this new benchmark dataset outperforms related approaches, achieving a board recognition accuracy of 15.26% ($\approx$7x better than the current state-of-the-art).
@@ 1 Introduction
Chess recognition from a single image poses a significant challenge in the field of computer vision. The task requires accurate identification of each chess piece’s type and position on a chessboard configuration. The predominant approach [30, 16, 6, 28] is to divide the overall task into the tasks of chessboard detection, square localization, and piece classification within each square. However, this approach not only suffers from error accumulation throughout the intermediate steps but also often sets restrictions to the task, with regard to selected viewing angles [28]. In this paper, we propose a novel approach for chess recognition that circumvents this pipeline and directly predicts the positions of the pieces, with respect to the chessboard, from the entire image.
One key advantage of our approach is that it does not impose any restrictions on the input image, such as specific angles and orientations. Our model can handle images taken from various perspectives, enabling robust chess recognition. By leveraging a deep neural network, the model is able to extract and use relevant visual features to efficiently predict the pieces’ type and positions. Additionally, our approach does not require any human input beyond the input image itself. Unlike traditional methods [22, 8, 17, 28] that rely on user input, such as manually selecting the corners of the chessboard or defining the player’s perspective, our approach directly learns to recognize the chessboard and infer the piece positions strictly from the visual information in the input image.
Furthermore, a factor that has hindered research on this particular domain is the lack of a comprehensive real-world dataset. To this end, we introduce a new real-world dataset specifically designed for chess recognition, Chess Recognition Dataset (ChessReD) [14], consisting of $10,800$ images and their corresponding annotations, that would allow us to effectively train and evaluate our proposed approach. The dataset comprises a diverse collection of chessboard images, covering various viewing angles, lighting conditions, camera specifications, and piece configurations. By providing this dedicated dataset, we enable both further research in chess recognition algorithms and a common benchmark for researchers to evaluate their approaches.
The results indicate that our approach in chess recognition on this new benchmark dataset outperforms related approaches, achieving a board recognition accuracy of 15.26% ($\approx$7x better than the current state-of-the-art). Thus, our contributions can be summarized as follows.
  • We introduce the first dataset of real images for chess recognition, with increased variability in terms of viewing angles and chess formations compared to synthetic alternatives
  • We demonstrate that chess recognition can be tackled using an end-to-end approach, with improved performance compared to sequential alternatives
@@ 2 Related Work
2.1 Early Approaches in Chess Recognition
@@
Chess recognition has been a subject of research in the field of computer vision, with several approaches proposed to tackle the challenges associated mainly with detecting the chessboard, but also with recognizing the individual pieces on top of it. Early attempts in the field primarily focused on integrating the chess recognition task as a part of a chess-playing robotic system [24, 22, 15, 1]. These systems detected chess moves by comparing the previous frame, with a known game state, to the current frame. They relied on detecting the occupied squares of the chessboard along with the colors of the pieces occupying them. As such, these methods were only able to detect valid chess moves and failed to detect events when two pieces of the same color were swapped, either illegally, or by promoting pawns to another piece type. Additionally, they worked under the assumption that the previous inferred state is correct. Thus, in case of an erroneous move prediction, all of the subsequent game states would be incorrect. Despite the aforementioned issues, the same approach has been adopted by several studies [25, 12, 3, 17, 4, 11] since, with consecutive frames obtained also from a video stream [22, 25, 12]. In this study, contrary to these approaches, we aim to develop a robust method that does not rely on the correctness of the previous inferred state but rather performs chess recognition from a single input image.
2.2 Chessboard Detection
@@
For cases when the previous state is unknown, chess recognition from a single image has also been the focus of studies. Same as with the previously mentioned approaches, the first step is to employ image processing techniques to detect the chessboard and the individual squares; a challenging task even on its own. While it can be simplified by explicitly asking the user to select the four corner points [22, 8, 17], modifying the chessboard [24, 1, 7] (e.g. using a reference frame around the chessboard), or setting constraints on the camera view angle (e.g. top-view) [24, 22, 1, 25, 12, 17, 16], Neufeld et al. [18] recognized that these approaches do not represent a general solution, where the chessboard could be in arbitrary locations or the image taken from various camera angles. They proposed a line-based detection method which they combined with probabilistic reasoning. However, they also restricted the setting by expecting the camera angle to be in the range of a human player’s perspective. Other studies have also exploited specific viewing angles, such as the players’ perspectives [8, 3, 4, 28] or side views [7, 19]. While chessboard detection utilizing the Harris corner detection algorithm [1, 12, 11], template matching [24, 15], or flood fill [25] have been explored, in accordance with [18], line-based chessboard detection methods have received significant research attention [3, 4, 30, 23, 31, 7, 16]. Czyzewski et al. [6] introduced an approach based on iterative heat map generation which visualizes the probability of a chessboard being located in a sub-region of the image. After each iteration, the four-sided area of the image containing the highest probability values is cropped and the process is repeated until convergence. While this method involves a great computational overhead, it is able to detect chessboards from images taken from varied angles, with poor quality, and regardless of the state of the actual chessboard (e.g. damaged chessboard with deformed edges), with a 99.6% detection accuracy. Wölflein and Arandjelović [28] proposed a chessboard detection method that leveraging the geometric nature of the chessboard, utilizes a RANSAC-based algorithm to iteratively refine the homography matrix and include all the computed intersection points. Their method demonstrated impressive results, since it successfully detected all of the chessboards in their validation dataset. However, it’s worth noting that the dataset only included images with viewing angles within the range of a player’s perspective. In our paper, we bypass the chessboard detection task, allowing the deep learning models to internally infer its position, and thus we do not rely on user input, or specific viewing angles.
2.3 Piece Classification
@@
Upon detection of the chessboard, the next step the aforementioned approaches employ is piece classification. A number of techniques have been developed to address this task, either in a 2-way approach (i.e. color and type), or 1-way by treating each combination of piece color and type as a separate class (e.g. “white-rook”). In Matuszek et al. [15], the authors utilized one classifier to determine the piece color, and then for each color they trained a type classifier using concatenated scale-invariant feature transform (SIFT) and kernel descriptors for features. A similar approach was used in Ding [8], where the author employed SIFT and histogram of oriented gradients (HOG) as feature descriptors for piece type classification with support vector machine (SVM) classifiers. The color was subsequently detected by comparing the binarized image of the square with that of an empty one. Danner and Kafafy [7] and Xie et al. [30] argued that the lack of distinguishable texture in small objects, such as chess pieces, leads to insufficient features obtained using SIFT descriptors. Both studies suggested a template matching approach for piece classification, comparing the contours of the detected pieces with reference templates obtained from various angles. Wei et al. [26] proposed an approach to recognize pieces using a depth camera and a volumetric convolutional neural network (CNN). More recent studies [16, 6, 19, 28] follow the 1-way approach for piece classification. They train CNNs to distinguish between 12 or 13 classes of objects (i.e. six piece types in both colors and one for empty squares), obtaining impressive results. In Czyzewski et al. [6], they also leverage domain knowledge, to improve piece classification, by utilizing a chess engine to calculate the most probable piece configurations and clustering similar figures into groups to deduce formations based on cardinalities. Additionally, given the variation in appearance between chess sets, Wölflein and Arandjelović [28] proposed a novel fine-tuning process for their piece classifier to unseen chess sets. In our paper, same as with the chessboard detection task, the classification of the pieces is performed by the deep learning model, without the need to train a separate piece classifier.
2.4 Chess Datasets
@@
A common problem frequently mentioned in literature [8, 16, 6, 28] is the lack of a comprehensive chess dataset. This issue hinders not only the ability to fairly evaluate the proposed methods in a common setting but also impedes the deployment of deep learning end-to-end approaches that require a vast amount of data. One proposed solution to this problem is the use of synthetic generated data. In [26], the authors produce point cloud data using a 3D computer-aided design (CAD) model, while Blender [5] was used to produce synthetic image datasets from a top view camera angle [17], or the player’s perspective [28]. In our paper, we introduce the first chess recognition dataset of real images, without setting any of the aforementioned restrictions regarding the viewing angles.
@@ 3 Chess Recognition Dataset (ChessReD)
The availability of large-scale annotated datasets is critical to the advancement of computer vision research. In this section, we tackle a main issue in the field of chess recognition (i.e. the lack of a comprehensive dataset) by presenting a novel dataset [14] specifically designed for this task. The dataset comprises a diverse collection of images of chess formations captured using smartphone cameras; a sensor choice made to ensure real-world applicability.
3.1 Data Collection and Annotation
The dataset was collected by capturing images of chessboards with various chess piece configurations. To guarantee the variability of those configurations, we relied upon the chess opening theory. The Encyclopedia of Chess Openings (ECO) classifies opening sequences into five volumes with 100 subcategories each that are uniquely identified by an ECO code. We randomly selected 20 ECO codes from each volume. Subsequently, each code of this set was randomly matched to an already played chess game that followed the particular opening sequence denoted by the ECO code; thus creating a set of 100 chess games. Finally, using the move-by-move information provided by Portable Game Notations (PGNs) that are used to record chess games, the selected games were played out on a physical chessboard with images being captured after each move.
Three distinct smartphone models were used to capture the images. Each model has different camera specifications, such as resolution and sensor type, that introduce further variability in the dataset. The images were also taken from diverse angles, ranging from top-view to oblique angles, and from different perspectives (e.g. white player perspective, side view, etc.). These conditions simulate real-world scenarios where chessboards can be captured from a bystander’s arbitrary point of view. Additionally, the dataset includes images captured under different lighting conditions, with both natural and artificial light sources introducing these variations. Most of these variations are illustrated in the four image samples of Figure 1. Each of those samples highlights a different challenge in chess recognition. Occlusions between pieces occur more often in images captured from a low angle (Fig. 0(c)) or a player’s perspective (Fig. 0(b)), while pieces are rarely occluded in top-view images (Fig. 0(d)). However, distinct characteristics of pieces (e.g. the queen’s crown) that could aid the chess recognition task are less distinguishable in a top-view.
@@ 3.1 Data Collection and Annotation
Figure 1: Image samples from ChessReD
3.2 Data Statistics
In this section, we present an overview of the ChessReD statistics. As mentioned in Section 3.1, the dataset consists of 100 chess games, each with an arbitrary number of moves and therefore images, amounting to a total of 10,800 images being collected. The dataset was split into training, validation, and test sets following an 60/20/20 split. Since two consecutive images of a chess game differ only by one move, the split was performed on game-level to ensure that quite similar images would not end up in different sets. The split was also stratified over the three distinct smartphone cameras (Apple iPhone 12, Huawei P40 pro, Samsung Galaxy S8) that were used to capture the images. Table 1 presents an overview of the image statistics per smartphone. The three smartphone cameras introduced variations to the dataset based on the distinct characteristics of their sensors. For instance, while the image resolution for the Huawei phone was 3072x3072, the resolution for the remaining two models was 3024x3024.
@@ 3.2 Data Statistics
Table 1: Overview of the image statistics
Table 2 presents an overview of the annotations in ChessReD. In LABEL:tab:stats-anns-pos we can see that there is a significant imbalance between provided annotations for the piece type “Pawn” and the rest of the types in the dataset. This was to be expected since every chess game starts with 8 pawns in each side and only one or two of the remaining piece types. Regarding the colors of the pieces, no imbalance is detected in the dataset. Additionally, while annotations about the position of the pieces in algebraic notation are available for every image in the dataset, we provide bounding box and chessboard corner annotations only for a subset of 20 randomly selected games from the train, validation, and test sets. For this subset we followed a 70/15/15 split stratified over the smartphone cameras, which led to a total of 14 training games (1442 images), 3 validation games (330 images), and 3 test games (306 images) being annotated. In LABEL:tab:stats-anns-bboxes we can see an overview of the annotation statistics for this subset, named ChessReD2K.
Figure 2: Bounding box and corner point annotations in ChessReD2K
@@
Table 2: Overview of the annotation statistics
4 End-to-End Chess Recognition
Unlike the conventional pipeline in chess recognition that involves chessboard detection, square localization, and chess piece classification, the focus of this study was to develop an end-to-end approach that tackles the recognition task utilizing only a single image as input. Thus, the developed method should take as input an image of a chessboard and output the type and the positions of the pieces relative to the board. To this end, we experimented with two different solutions by treating the problem either as a multi-class multi-label classification or as a relative object detection task.
4.1 Classification Approach
In the classification approach, we treat each chessboard square as a distinct label. Since there are 64 squares in each image, and thus 64 labels, this problem is treated as a multi-label classification task. Additionally, each square in the chessboard is either unoccupied or occupied by one of the 12 different types of pieces (i.e. 6 per color) in chess. Therefore, to each label we assign one of 13 classes (i.e. 12 piece types and empty), constituting this approach multi-label multi-class classification. By formulating this approach as such, the goal is for the model to learn the intricate relationships and visual patterns associated with the individual squares.
4.2 Relative Object Detection Approach

### v3

@@ Abstract
Chess recognition is the task of extracting the chess piece configuration from a chessboard image. Current approaches use a pipeline of separate, independent, modules such as chessboard detection, square localization, and piece classification. Instead, we follow the deep learning philosophy and explore an end-to-end approach to directly predict the configuration from the image, thus avoiding the error accumulation of the sequential approaches and eliminating the need for intermediate annotations. Furthermore, we introduce a new dataset, Chess Recognition Dataset (ChessReD), that consists of 10,800 real photographs and their corresponding annotations. In contrast to existing datasets that are synthetically rendered and have only limited angles, ChessReD has photographs captured from various angles using smartphone cameras; a sensor choice made to ensure real-world applicability. Our approach in chess recognition on the introduced challenging benchmark dataset outperforms related approaches, successfully recognizing the chess pieces’ configuration in 15.26% of ChessReD’s test images. This accuracy may seem low, but it is $\approx$7x better than the current state-of-the-art and reflects the difficulty of the problem. The code and data are available through: https://github.com/ThanosM97/end-to-end-chess-recognition.
Refer to caption Refer to caption
Figure 1: Chess recognition input image and output configuration.
@@ 1 INTRODUCTION
Parsing a chess position from a (smartphone) image eases match analysis and facilitates coaching of children, offering chess positions without writing each move down (Figure 1). Chess position recognition requires accurate identification of each chess piece’s type and position on a chessboard configuration. The predominant approach [Xie et al., 2018a, Mehta, 2020, Czyzewski et al., 2020, Wölflein and Arandjelović, 2021] is to split it into independent sequential sub-tasks: chessboard detection, square localization, and piece classification within each square. We here depart from the observation that such independent sub-tasks suffer from error accumulation throughout each intermediate independent step. In this paper, we do away with the sequential independent sub-tasks and propose an end-to-end approach for chess recognition, directly predicting the positions of the pieces, with respect to the chessboard, from the entire image.
One key advantage of our approach is that it does not require any human input beyond the input image itself, unlike traditional methods [Sokic and Ahic-Djokic, 2008, Ding, 2016, Neto and Campello, 2019, Wölflein and Arandjelović, 2021] that rely on user input, such as manually selecting the corners of the chessboard or defining the player’s perspective. By leveraging a deep neural network, the model is able to extract and use relevant visual features to efficiently recognize the chessboard, predict the pieces’ type and positions directly from a single image.
To evaluate and facilitate our learning-based approach we need a large and good quality dataset. Currently, there is no real-world chess recognition dataset and thus, we introduce our own novel real-world Chess Recognition Dataset (ChessReD) , which we openly share with the community [^1] https://data.4tu.nl/datasets/99b5c721-280b-450b-b058-b2900b69a90f . Our ChessReD dataset consists of $10,800$ images and their corresponding annotations, allowing us to effectively train and evaluate approaches. Although there might be a human preference for certain viewpoint [Van Gemert, 2011], we capture a diverse collection of chessboard images, covering various viewing angles, lighting conditions, camera specifications, and piece configurations. This dataset enables both further research in chess recognition algorithms and a realistic benchmark.
Our approach in chess recognition on our ChessReD benchmark outperforms related approaches, achieving a board recognition accuracy of 15.26% ($\approx$7x better than the current state-of-the-art). Our contributions can be summarized as follows.
  • We introduce ChessReD: the first dataset of real images for chess recognition, with high viewing angle variability and diversity in chess formations, compared to synthetic alternatives.
  • We demonstrate end-to-end chess recognition with improved performance compared to alternatives that rely on a sequential pipeline of independent components.
@@ 2 RELATED WORK
Early Approaches in Chess Recognition
@@
Chess recognition has been a subject of research in the field of computer vision, with several approaches proposed to tackle the challenges associated mainly with detecting the chessboard, but also with recognizing the individual pieces on top of it. Early attempts in the field primarily focused on integrating the chess recognition task as a part of a chess-playing robotic system [Urting and Berbers, 2003, Sokic and Ahic-Djokic, 2008, Matuszek et al., 2011, Banerjee et al., 2011]. These systems detected chess moves by comparing the previous frame, with a known game state, to the current frame. They relied on detecting the occupied squares of the chessboard along with the colors of the pieces occupying them. As such, these methods were only able to detect valid chess moves and failed to detect events when two pieces of the same color were swapped, either illegally, or by promoting pawns to another piece type. Additionally, they worked under the assumption that the previous inferred state is correct. Thus, in case of an erroneous move prediction, all of the subsequent game states would be incorrect. Despite the aforementioned issues, the same approach has been adopted by several studies [Wang and Green, 2013, Koray et al., 2016, Chen et al., 2016, Neto and Campello, 2019, Chen and Wang, 2019, Kołosowski et al., 2020] since, with consecutive frames obtained also from a video stream [Sokic and Ahic-Djokic, 2008, Wang and Green, 2013, Koray et al., 2016]. In our paper, contrary to these approaches, we aim to develop a robust method that does not rely on the correctness of the previous inferred state but rather performs chess recognition from a single input image.
Chessboard Detection
@@
For cases when the previous state is unknown, chess recognition from a single image has also been the focus of studies. Same as with the previously mentioned approaches, the first step is to employ image processing techniques to detect the chessboard and the individual squares; a challenging task even on its own. While it can be simplified by explicitly asking the user to select the four corner points [Sokic and Ahic-Djokic, 2008, Ding, 2016, Neto and Campello, 2019], modifying the chessboard [Urting and Berbers, 2003, Banerjee et al., 2011, Danner and Kafafy, 2015] (e.g. using a reference frame around the chessboard), or setting constraints on the camera view angle (e.g. top-view) [Urting and Berbers, 2003, Sokic and Ahic-Djokic, 2008, Banerjee et al., 2011, Wang and Green, 2013, Koray et al., 2016, Neto and Campello, 2019, Mehta, 2020], Neufeld et al. [Neufeld and Hall, 2010] recognized that these approaches do not represent a general solution, where the chessboard could be in arbitrary locations or the image taken from various camera angles. They proposed a line-based detection method which they combined with probabilistic reasoning. However, they also restricted the setting by expecting the camera angle to be in the range of a human player’s perspective. Other studies have also exploited specific viewing angles, such as the players’ perspectives [Ding, 2016, Chen et al., 2016, Chen and Wang, 2019, Wölflein and Arandjelović, 2021] or side views [Danner and Kafafy, 2015, Quintana et al., 2020]. While chessboard detection utilizing the Harris corner detection algorithm [Banerjee et al., 2011, Koray et al., 2016, Kołosowski et al., 2020], template matching [Urting and Berbers, 2003, Matuszek et al., 2011], or flood fill [Wang and Green, 2013] have been explored, in accordance with [Neufeld and Hall, 2010], line-based chessboard detection methods have received significant research attention [Chen et al., 2016, Chen and Wang, 2019, Xie et al., 2018a, Tam et al., 2008, Xie et al., 2018b, Danner and Kafafy, 2015, Mehta, 2020]. Czyzewski et al. [Czyzewski et al., 2020] introduced an approach based on iterative heat map generation which visualizes the probability of a chessboard being located in a sub-region of the image. After each iteration, the four-sided area of the image containing the highest probability values is cropped and the process is repeated until convergence. While this method involves a great computational overhead, it is able to detect chessboards from images taken from varied angles, with poor quality, and regardless of the state of the actual chessboard (e.g. damaged chessboard with deformed edges), with a 99.6% detection accuracy. Wölflein and Arandjelović [Wölflein and Arandjelović, 2021] proposed a chessboard detection method that leveraging the geometric nature of the chessboard, utilizes a RANSAC-based algorithm to iteratively refine the homography matrix and include all the computed intersection points. Their method demonstrated impressive results, since it successfully detected all of the chessboards in their validation dataset. However, it’s worth noting that the dataset only included images with viewing angles within the range of a player’s perspective. In our paper, we bypass the chessboard detection task, allowing the deep learning models to internally infer its position, and thus we do not rely on user input, or specific viewing angles.
Piece Classification
@@
Upon detection of the chessboard, the next step the aforementioned approaches employ is piece classification. A number of techniques have been developed to address this task, either in a 2-way approach (i.e. color and type), or 1-way by treating each combination of piece color and type as a separate class (e.g. “white-rook”). In Matuszek et al. [Matuszek et al., 2011], the authors utilized one classifier to determine the piece color, and then for each color they trained a type classifier using concatenated scale-invariant feature transform (SIFT) and kernel descriptors for features. A similar approach was used in Ding [Ding, 2016], where the author employed SIFT and histogram of oriented gradients (HOG) as feature descriptors for piece type classification with support vector machine (SVM) classifiers. The color was subsequently detected by comparing the binarized image of the square with that of an empty one. Danner and Kafafy [Danner and Kafafy, 2015] and Xie et al. [Xie et al., 2018a] argued that the lack of distinguishable texture in small objects, such as chess pieces, leads to insufficient features obtained using SIFT descriptors. Both studies suggested a template matching approach for piece classification, comparing the contours of the detected pieces with reference templates obtained from various angles. Wei et al. [Wei et al., 2017] proposed an approach to recognize pieces using a depth camera and a volumetric convolutional neural network (CNN). More recent studies [Mehta, 2020, Czyzewski et al., 2020, Quintana et al., 2020, Wölflein and Arandjelović, 2021] follow the 1-way approach for piece classification. They train CNNs to distinguish between 12 or 13 classes of objects (i.e. six piece types in both colors and one for empty squares), obtaining impressive results. In Czyzewski et al. [Czyzewski et al., 2020], they also leverage domain knowledge, to improve piece classification, by utilizing a chess engine to calculate the most probable piece configurations and clustering similar figures into groups to deduce formations based on cardinalities. Additionally, given the variation in appearance between chess sets, Wölflein and Arandjelović [Wölflein and Arandjelović, 2021] proposed a novel fine-tuning process for their piece classifier to unseen chess sets. In our paper, same as with the chessboard detection task, the classification of the pieces is performed by the deep learning model, without the need to train a separate piece classifier.
Chess Datasets
@@
A common problem frequently mentioned in literature [Ding, 2016, Mehta, 2020, Czyzewski et al., 2020, Wölflein and Arandjelović, 2021] is the lack of a comprehensive chess dataset. This issue hinders not only the ability to fairly evaluate the proposed methods in a common setting but also impedes the deployment of deep learning end-to-end approaches that require a vast amount of data. One proposed solution to this problem is the use of synthetic generated data. In [Wei et al., 2017], the authors produce point cloud data using a 3D computer-aided design (CAD) model, while Blender [Community, 2018] was used to produce synthetic image datasets from a top view camera angle [Neto and Campello, 2019], or the player’s perspective [Wölflein and Arandjelović, 2021]. In our paper, we introduce the first chess recognition dataset of real images, without setting any of the aforementioned restrictions regarding the viewing angles.
@@ 3 CHESS RECOGNITION DATASET (ChessReD)
The availability of large-scale annotated datasets is critical to the advancement of computer vision research. In this section, we tackle a main issue in the field of chess recognition (i.e. the lack of a comprehensive dataset) by presenting a novel dataset1 specifically designed for this task. The dataset comprises a diverse collection of photographs of chess formations captured using smartphone cameras; a sensor choice made to ensure real-world applicability.
Data Collection and Annotation
The dataset was collected by capturing photographs of chessboards with various chess piece configurations. To guarantee the variability of those configurations, we relied upon the chess opening theory. The Encyclopedia of Chess Openings (ECO) classifies opening sequences into five volumes with 100 subcategories each that are uniquely identified by an ECO code. We randomly selected 20 ECO codes from each volume. Subsequently, each code of this set was randomly matched to an already played chess game that followed the particular opening sequence denoted by the ECO code; thus creating a set of 100 chess games. Finally, using the move-by-move information provided by Portable Game Notations (PGNs) that are used to record chess games, the selected games were played out on a single physical chessboard, commonly used at chess clubs, with images being captured after each move.
Three distinct smartphone models were used to capture the images. Each model has different camera specifications, such as resolution and sensor type, that introduce further variability in the dataset. The images were also taken from diverse angles, ranging from top-view to oblique angles, and from different perspectives (e.g. white player perspective, side view, etc.). These conditions simulate real-world scenarios where chessboards can be captured from a bystander’s arbitrary point of view. Additionally, the dataset includes images captured under different lighting conditions, with both natural and artificial light sources introducing these variations. Most of these variations are illustrated in the four image samples of Figure 2. Each of those samples highlights a different challenge in chess recognition. Occlusions between pieces occur more often in images captured from a low angle (Figure 1(c)) or a player’s perspective (Figure 1(b)), while pieces are rarely occluded in top-view images (Figure 1(d)). However, distinct characteristics of pieces (e.g. the queen’s crown) that could aid the chess recognition task are less distinguishable in a top-view. More samples of ChessReD can be seen in the Appendix.
@@
Figure 2: Image samples from ChessReD.
Data Statistics
The dataset consists of 100 chess games, each with an arbitrary number of moves and therefore images, amounting to a total of 10,800 images being collected. The dataset was split into training, validation, and test sets following an 60/20/20 split. Since two consecutive images of a chess game differ only by one move, the split was performed on game-level to ensure that quite similar images would not end up in different sets. The split was also stratified over the three distinct smartphone cameras (Apple iPhone 12, Huawei P40 pro, Samsung Galaxy S8) that were used to capture the images. Table 1 presents an overview of the image statistics per smartphone. The three smartphone cameras introduced variations to the dataset based on the distinct characteristics of their sensors. For instance, while the image resolution for the Huawei phone was 3072x3072, the resolution for the remaining two models was 3024x3024.
Table 1: Overview of the image statistics.
@@
Table 2 presents an overview of the annotations in ChessReD. In Table 2 illustrates a significant imbalance between annotations for the piece type “Pawn” and the rest of the pieces. This was to be expected since every chess game starts with 8 pawns in each side and only one or two of the remaining piece types. Regarding the colors of the pieces, no imbalance is detected in the dataset. Additionally, while annotations about the position of the pieces in algebraic notation are available for every image in the dataset, we provide bounding box and chessboard corner annotations only for a subset of 20 randomly selected games (2,078 images) from the train, validation, and test sets. For this subset we followed a 70/15/15 split stratified over the smartphone cameras, which led to a total of 14 training games (1442 images), 3 validation games (330 images), and 3 test games (306 images) being annotated. In Table 2 we can see an overview of the annotation statistics for this subset, named ChessReD2K.
Figure 3: Bounding box and corner point annotations in ChessReD2K.
Table 2: Overview of the annotation statistics.
(a) Piece positions on the chessboard in ChessReD
(b) Bounding boxes in ChessReD2K
@@
(b) Bounding boxes in ChessReD2K
4 METHOD: END-TO-END CHESS RECOGNITION
Unlike the conventional pipeline in chess recognition that involves separate, independent, modules of chessboard detection, square localization, and chess piece classification, the focus of this study was to explore an end-to-end approach that tackles the recognition task utilizing only a single image as input. Thus, the developed method should take as input an image of a chessboard and output the type and the positions of the pieces relative to the board. To this end, we experimented with two different solutions by treating the problem either as a multi-class multi-label classification or as a relative object detection task.
Classification approach
Here, each chessboard square is a distinct label. Since there are 64 squares in each image, and thus 64 labels, this problem is treated as a multi-label classification task. Each square in the chessboard is either unoccupied or occupied by one of the 12 different types of pieces (i.e. 6 per color) in chess. Therefore, to each label we assign one of 13 classes (i.e. 12 piece types and ’empty’). By formulating it as multi-label multi-class classification, the goal is for the model to learn the intricate relationships and visual patterns associated with the individual squares.
Relative Object Detection Approach